# VideoQA-SC: Adaptive Semantic Communication for Video Question Answering

### v1

@@ 2.6 Training Strategy
2.6.1 VideoQA Model Training
@@
2.6.2 Fixed-Bandwidth DJSCC Transmission Training
@@
2.6.3 Bandwidth-Adaptive DJSCC Transmission Training
@@
$\mathcal{L}_{\text{stage3}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{rate}}=\operatorname{CE}(\mathbf{a}_{\text{pred}},\mathbf{a}_{\text{label}})+\sum_{i}\sum_{j}\mathbf{M}_{i,j},$ (30)
@@
2.6.4 E2E Finetuning
Unfreezing VideoQA model parameters $\boldsymbol{\zeta}$ and $\boldsymbol{\nu}$. Jointly finetuning all system parameters $\boldsymbol{\zeta}$, $\boldsymbol{\theta}$, $\boldsymbol{\epsilon}$, $\boldsymbol{\phi}$ and $\boldsymbol{\nu}$ to improve the E2E system performance. This stage has the same loss as $\mathcal{L}_{\text{stage3}}$:
$\mathcal{L}_{\text{stage4}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{rate}}=\operatorname{CE}(\mathbf{a}_{\text{pred}},\mathbf{a}_{\text{label}})+\sum_{i}\sum_{j}\mathbf{M}_{i,j}.$ (31)
@@ 2.6.4 E2E Finetuning
3 Experiments
In this section, we introduce the experimental setup and provide quantified experimental results to demonstrate the effectiveness of VideoQA-SC for performing VideoQA tasks. We compare VideoQA-SC with SC systems adopted traditional SSCC and advanced DJSCC transmission schemes under various channel conditions and bandwidth constraints.
3.1 Experimental Setup
3.1.1 Datasets
@@
3.1.2 Implementation Details
In our experiments, $l_{v}$, ${l_{c}}$, ${l_{f}}$ are set to $16$, $4$, $4$ for sparse sampling. $r$ and $m$ are set to $10$ and $2048$ for spatiotemporal semantic encoder. $d$ is set to $256$ for DJSCC transmission, which also represents the maximum bandwidth that can be allocated for each token. Both the JSC encoder and the JSC decoder consist of $L=4$ Transformer blocks for JSC encoding/decoding. There are also $L=4$ rate predictors for bandwidth allocation during JSC encoding. $q$ is set to $8$ for adaptive bandwidth allocation, which denotes the set of candidate retained channels for each token is =$\mathcal{R}=\left\{2,4,8,16,32,64,128,256\right\}$. Each question has $b=5$ candidate answers to form QA pairs.
@@ 3.1.2 Implementation Details
3.1.3 Comparison Schemes
We compare the VideoQA-SC with traditional SSCC-based SC systems and DJSCC-based SC systems that perform VideoQA based on reconstructed videos. Specifically, we use SSCC and DJSCC schemes to transmit videos over the noisy wireless channel, and use the same optimized VideoQA model($g(\cdot;\boldsymbol{\zeta})$ and $u(\cdot;\boldsymbol{\nu})$) to perform VideoQA tasks.
@@ 3.2 Ablation Study
3.3 Performance comparison of different SC models
As depicted in Fig. 8 and 9, the proposed VideoQA-SC is compared with SSCC-based and DJSCC-based SC systems over the AWGN channel under different BCR constraints and SNRs. We use the optimized VideoQA model ($g(\cdot;\boldsymbol{\zeta})$ and $u(\cdot;\boldsymbol{\nu})$) with lossless video and text to execute VideoQA tasks as the upper bound performance of the VideoQA-SC. The lower bound performance of the VideoQA-SC is obtained by performing VideoQA tasks with only text information. For fair comparison, the same VideoQA model optimized in training stage 1 is adopted in all comparison SC schemes. Content-adaptive and SNR-adaptive VideoQA-SCs are trained and tested under fixed SNRs and mixed SNRs, respectively.
@@ 3.3 Performance comparison of different SC models
By comparing Fig. 8LABEL:sub@Fig:A_0dB, LABEL:sub@Fig:A_5dB and LABEL:sub@Fig:A_10dB or Fig. 9LABEL:sub@Fig:T_0dB, LABEL:sub@Fig:T_5dB and LABEL:sub@Fig:T_10dB, it can be found that the proposed VideoQA-SC outperforms other SC schemes especially in low SNRs (00 dB). Since comparison SC systems need to restore the original video at the pixel level, they all have a turning point with the necessary transmission bandwidth and the corresponding minimum BCR $B_{\text{min}}$. When the BCR constraint $B<B_{\text{min}}$, comparison SC systems lack the necessary information to reconstruct the original videos, which leads to a rapid deterioration of VideoQA performance. As the SNR decreases, $B_{\text{min}}$ gradually increases, which indicates that more bandwidths are required to guarantee the basic performance of SC systems when the channel conditions are worse. The DJSCC-based video transmission model DVST also employs a bandwidth-adaptive approach and achieves excellent video reconstruction performance, e.g., peak signal-to-noise ratio with the same bandwidth. However, there are slight bandwidth savings for DJSCC-based SC systems compared with SSCC-based SC systems for VideoQA tasks. It indicates that the adopted entropy-based bandwidth allocation limited to pixel-level video reconstruction is not effective for SC systems oriented to a particular intelligent task. Besides, although DJSCC-based SC systems can avoid the cliff effect and are robust to channel noise, they still face the risk of system collapse when $B<B_{\text{min}}$ cannot be satisfied.
@@
Different from comparison schemes at pixel-level, the proposed VideoQA-SC extracts and processes video semantics at the object and frame levels, and directly performs VideoQA tasks based on the reconstructed video semantics at the receiver, which results in significant bandwidth savings with guaranteed VideoQA performance. In this way, the minimum BCR $B_{\text{min}}$ of VideoQA-SC can be very low, which ensures the basic performance of VideoQA-SC under a wide range of BCR constraints. It can be seen that both types of VideoQA-SC achieve accuracy of over $53\%$ and $64\%$ under all testing SNR and BCR constraints in the TGIF-QA-R Action and Transition dataset, respectively. Benefiting from E2E joint training and SNR-adaptive bandwidth allocation, the proposed VideoQA-SC demonstrates robustness to noisy wireless channels while approaching the upper bound performance. In Fig. 8LABEL:sub@Fig:A_0dB, the accuracy of SNR-adaptive VideoQA-SC exceeds that of the DJSCC-based SC system by $5.17\%$ with almost $\frac{1}{200}$ of its bandwidth.
@@
We further show the performance of VideoQA-SC over Rayleigh block fading channels with different fading parameters $\sigma_{h}$. For each $\hat{\mathbf{s}}_{v}^{c}$, a corresponding $h$ is sampled from the Rayleigh distribution $\operatorname{Rayleigh}(\sigma_{h})$. We adjust the signal power so that the ratio of it to the simulated average noise power is $1$. Then, the statistical SNR and the current SNR are fed into rate predictors to improve the robustness of SNR-adaptive VideoQA-SC to Rayleigh block fading channels. Fig. 10 shows that VideoQA-SC is able to achieve slow and smooth degradation of VideoQA performance with the increasing fading, however, DJSCC-based SC fails to resist noise in Rayleigh block fading channels.
@@
Fig. 11 provides a visual example to show the robustness of VideoQA-SC to the AWGN channel. DVST is adopted as the comparison scheme for video transmission. In this case, every $4$ frames are combined into one clip for semantic extraction and video coding. The first frame is used as I-frame to give the reference of the coding of P-frames for DVST.
@@
Figure 11: A visualization of VideoQA-SC robust to the channel noise. Each row represents the bandwidth allocation made by the corresponding model for all video frames over the AWGN channel with a specific SNR. Each column denotes the number of complex transmission symbols allocated for the corresponding frame in different cases. The correct answer is marked in red.
4 Conclusion
In this paper, we have proposed an E2E multimodal SC system, VideoQA-SC, to perform VideoQA tasks in wireless networks without relying on the reconstructed original source. Taking advantage of the efficient video semantic extraction and the learning-based bandwidth-adaptive DJSCC transmission, VideoQA-SC is able to fully leverage video semantic information to improve system performance for VideoQA tasks with high bandwidth-efficiency and noise-robustness. VideoQA-SC is trained in an E2E manner with the goal of maximize the VideoQA performance under bandwidth constraints. Experiments show that the proposed VideoQA-SC outperforms traditional SSCC-based and advanced DJSCC-based SC systems under a wide range of channel conditions and bandwidth constraints. In particular, VideoQA-SC improves $5.17\%$ VideoQA accuracy while achieves nearly $99.5\%$ bandwidth savings compared with the DJSCC-based SC system over AWGN channel when SNR is 00 dB, which demonstrates the great potential of task-oriented SC system design for video applications.

### v2

@@ III-F Training Strategy
III-F1 VideoQA Model Training
@@
III-F2 Fixed-Bandwidth DJSCC Transmission Training
@@
III-F3 Bandwidth-Adaptive DJSCC Transmission Training
@@
$\mathcal{L}_{\text{stage3}}$ $=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{rate}}$ (30)
$=\operatorname{CE}(\mathbf{a}_{\text{pred}},\mathbf{a}_{\text{label}})+\lambda\sum_{i}\sum_{j}\mathbf{M}_{i,j},$
@@
III-F4 E2E Finetuning
Unfreeze VideoQA model parameters $\boldsymbol{\zeta}$ and $\boldsymbol{\nu}$. Jointly finetune all system parameters $\boldsymbol{\zeta}$, $\boldsymbol{\theta}$, $\boldsymbol{\epsilon}$, $\boldsymbol{\phi}$ and $\boldsymbol{\nu}$ to improve the E2E system performance. This stage has the same loss as $\mathcal{L}_{\text{stage3}}$:
$\mathcal{L}_{\text{stage4}}$ $=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{rate}}$ (31)
$=\operatorname{CE}(\mathbf{a}_{\text{pred}},\mathbf{a}_{\text{label}})+\lambda\sum_{i}\sum_{j}\mathbf{M}_{i,j}.$
@@ III-F4 E2E Finetuning
IV Experiments
In this section, we introduce the experimental setup and provide quantified experimental results to demonstrate the effectiveness of VideoQA-SC for performing VideoQA tasks. We compare VideoQA-SC with traditional systems based on separate source channel coding (SSCC) and SC systems adopted advanced DJSCC transmission schemes under various channel conditions and bandwidth constraints.
IV-A Experimental Setup
IV-A1 Datasets
@@
IV-A2 Implementation Details
In our experiments, $l_{v}$, ${l_{c}}$, ${l_{f}}$ are set to $16$, $4$, $4$ for sparse sampling. $r$ and $m$ are set to $10$ and $2048$ for spatiotemporal semantic encoder. Note that the pretrained object-level feature extractor and the pretrained frame-level feature extractor served as preprocessing networks are not involved in the training process. $r$ is set to $10$ means that we only select the top $10$ detected objects with the highest confidence for each frame, which is a trade-off between network complexity and network performance. When the number of detected objects is less than $10$, we pad the additional object features with zeros to ensure a total of $10$ detected objects. $d$ is set to $256$ for DJSCC transmission, which also represents the maximum bandwidth that can be allocated for each token. Both the JSC encoder and the JSC decoder consist of $L=4$ Transformer blocks for JSC encoding/decoding. There are also $L=4$ rate predictors for bandwidth allocation during JSC encoding. $q$ is set to $8$ for adaptive bandwidth allocation, which denotes the set of candidate retained channels for each token is =$\mathcal{R}=\left\{2,4,8,16,32,64,128,256\right\}$. Each question has $b=5$ candidate answers to form QA pairs.
@@ IV-A2 Implementation Details
IV-A3 Comparison Schemes
We compare the VideoQA-SC with SSCC-based traditional systems and DJSCC-based SC systems that perform VideoQA based on reconstructed videos. Specifically, we use SSCC and DJSCC schemes to transmit videos over the noisy wireless channel, and use the same optimized VideoQA model ($g(\cdot;\boldsymbol{\zeta})$ and $u(\cdot;\boldsymbol{\nu})$) to perform VideoQA tasks.
@@ IV-B Ablation Study
IV-C Performance Comparison of Different Models
As depicted in Fig. 8 and 9, the proposed VideoQA-SC is compared with SSCC-based traditional systems and the DJSCC-based SC system over the AWGN channel under different BCR constraints and SNRs. We use the optimized VideoQA model ($g(\cdot;\boldsymbol{\zeta})$ and $u(\cdot;\boldsymbol{\nu})$) with lossless video and text to execute VideoQA tasks as the upper bound performance of the VideoQA-SC. The lower bound performance of the VideoQA-SC is obtained by performing VideoQA tasks only using texts at the receiver without video transmission. For fair comparison, the same VideoQA model optimized in training stage 1 is adopted in all comparison SC schemes. Content-adaptive and SNR-adaptive VideoQA-SCs are trained and tested under fixed SNRs and mixed SNRs, respectively.
@@ IV-C Performance Comparison of Different Models
By comparing Fig. 8LABEL:sub@Fig:A_0dB, LABEL:sub@Fig:A_5dB and LABEL:sub@Fig:A_10dB or Fig. 9LABEL:sub@Fig:T_0dB, LABEL:sub@Fig:T_5dB and LABEL:sub@Fig:T_10dB, it can be found that the proposed VideoQA-SC outperforms other SC schemes especially in low SNRs (00 dB). Since comparison SC systems need to restore the original video at the pixel level, When the BCR constraint $R<R_{\text{min}}$, comparison SC systems lack the necessary information to reconstruct the original videos, which leads to a rapid deterioration of VideoQA performance. As the SNR decreases, $R_{\text{min}}$ gradually increases, which indicates that more bandwidths are required to guarantee the basic performance of SC systems when the channel conditions are worse. The DJSCC-based video transmission model DVST also employs a bandwidth-adaptive approach and achieves excellent video reconstruction performance, e.g., peak signal-to-noise ratio with the same bandwidth. However, there are slight bandwidth savings for DJSCC-based SC systems compared with SSCC-based traditional systems for VideoQA tasks. It indicates that the adopted entropy-based bandwidth allocation limited to pixel-level video reconstruction is not effective for SC systems oriented to a particular intelligent task. Besides, although DJSCC-based SC systems can avoid the cliff effect and are robust to channel noise, they still face the risk of system collapse when $R<R_{\text{min}}$ cannot be satisfied.
Figure 11: Visualization of spatiotemporal correlations extracted by the semantic encoder.
@@
Different from comparison schemes at pixel-level, the proposed VideoQA-SC extracts and processes video semantics at the object and frame levels, and directly performs VideoQA tasks based on the reconstructed video semantics at the receiver, which results in significant bandwidth savings with guaranteed VideoQA performance. In this way, the minimum BCR $R_{\text{min}}$ of VideoQA-SC can be very low, which ensures the basic performance of VideoQA-SC under a wide range of BCR constraints. It can be seen that both types of VideoQA-SC achieve accuracy of over $53\%$ and $64\%$ under all testing SNR and BCR constraints in the TGIF-QA-R Action and Transition dataset, respectively. Benefiting from E2E joint training and SNR-adaptive bandwidth allocation, the proposed VideoQA-SC demonstrates robustness to noisy wireless channels while approaching the upper bound performance. In Fig. 8LABEL:sub@Fig:A_0dB, the accuracy of SNR-adaptive VideoQA-SC exceeds that of the DJSCC-based SC system by $5.17\%$ with almost $0.05\%$ of its bandwidth.
@@
We further show the performance of VideoQA-SC over Rayleigh block fading channels with different fading parameters $\sigma_{h}$ and fixed $\sigma_{n}^{2}=1$. In order to fix $\sigma_{n}^{2}$, we adjust the signal power to control the ratio of it to the simulated average noise power is $1$. If the perfect channel state information can be obtained at the transmitter, the statistical SNR and the current SNR are fed into rate predictors to improve the robustness of SNR-adaptive VideoQA-SC to Rayleigh block fading channels. Fig. 10 shows that VideoQA-SC is able to achieve slow and smooth degradation of VideoQA performance with the increasing fading, however, DJSCC-based SC fails to resist noise in Rayleigh block fading channels.
Figure 12: Visualization of different initial values of rate embeddings. Each row denotes trained initial values of rate embeddings $\mathbf{Y}_{\text{rate}}$ under corresponding $\lambda$ (higher $\lambda$ corresponds to lower BCR $R$). The horizontal axis indicates the frame index and the initial values are normalized over all frames. $\sigma$ indicates the standard deviation of corresponding rate embeddings.
Figure 13: An example of VideoQA-SC robust to the channel noise. Each row represents the bandwidth allocation made by the corresponding model for all video frames over the AWGN channel with a specific SNR. Each column denotes the number of complex transmission symbols allocated for the corresponding frame in different cases. The correct answer is marked in red.
IV-D Visualization and Analysis
We visualize the extracted semantics $\mathbf{Y}_{v}$ and the rate embeddings $\mathbf{Y}_{\text{rate}}$ trained under different $\lambda$ values to further analyze the operational mechanism of VideoQA-SC. Additionally, we conduct further experiments to demonstrate the ability of VideoQA-SC to handle different types of questions.
As shown in Fig. 11, we first visualize the video semantics $\mathbf{Y}_{v}$ extracted by the semantic encoder to observe how it models the spatiotemporal correlations of videos. Given that the semantic encoder extract spatiotemporal correlations within one clip, we visualize one clip including four frames from each of two different videos. We employ the Grad-CAM [47] method to obtain heatmaps of $\mathbf{Y}_{v}$ by overlaying the gradients obtained from backpropagation onto the original frames.
By comparing the frames in the first video, we can observe that the pixels near the football and the man’s legs exhibit higher gradients, indicated by warmer colors. This demonstrates that the spatiotemporal extraction pay more attention to the interaction between the football and the man, specifically the action of “kicking the football”. Furthermore, the visualization of the second video illustrates the extraction of spatiotemporal correlations when additional objects suddenly appear in some frames. For the first two frames, the gradients are concentrated near the man’s hand and the wall, suggesting that the extraction focus on the action of “graffiti”. However, when a bystander appears, the extraction pay attention to the interaction between the bystander and the man gradually, as the specific question is unknown. This ensures that the extracted video semantics are as comprehensive as possible and can be used to answer various questions.
As shown in Fig. 12, we then visualize the values of the trained initial values of rate embeddings $\mathbf{Y}_{\text{rate}}$ under three $\lambda$ values to explore their effect on the model. Given that the video semantics $\mathbf{Y}_{v}$ include 16 frames in our implementation, the initial values of $\mathbf{Y}_{\text{rate}}$ are normalized over all 16 frames, indicating the attention allocation to different frames. We can observe that the initial values corresponding to $\lambda_{1}$ exhibit extreme values in specific frames, which can be reflected in a higher $\sigma$. It demonstrates that when the bandwidth is limited, the network tends to perform VideoQA by some specific key frames. As $\lambda$ decreases from $\lambda_{1}$ to $\lambda_{3}$ ($R$ increases accordingly), the attention allocated to all frames approaches uniformity, which can be reflected in a lower $\sigma$. It indicates that when the available bandwidth is abundant, the model tends to utilize all frames equally to perform VideoQA.
@@ IV-D Visualization and Analysis
Fig. 13 provides an example to show the robustness of VideoQA-SC to the AWGN channel. DVST is adopted as the comparison scheme for video transmission. In this case, every $4$ frames are combined into one clip for semantic extraction and video coding. The first frame is used as I-frame to give the reference of the coding of P-frames for DVST.
@@
TABLE I: Performance of VideoQA-SC on NExT-QA dataset.
Model Causal Temporal Descriptive Average
Unlimited Bandwidth 50.17% 53.03% 61.64% 52.83%
$\lambda=0.001$ Accuracy -0.35% -2.82% -2.78% -1.41%
BCR 59.08 59.12 69.00 60.72
$\lambda=0.0015$ Accuracy -1.10% -3.01% -2.92% -1.88%
BCR 24.00 24.68 29.04 25.04
$\lambda=0.002$ Accuracy -1.37% -3.12% -4.49% -2.32%
BCR 7.44 7.61 8.76 7.71
  • 1
    For models trained with bandwidth constraints, average BCR are normalized by $10^{-6}$.
  • 2
    The training and testing are conducted under $\text{SNR}=10\text{dB}$.
We further test VideoQA-SC on NExT-QA[48] dataset to provide a more detailed analysis of VideoQA tasks. Compared with TGIF-QA-R dataset, NExT-QA dataset including various question types (causal, temporal and descriptive) with higher video resolutions and longer durations. The results are shown in TABLE. I. We use the trained VideoQA model with lossless video transmission (unlimited bandwidth) as the upper bound of comparison to measure the performance of VideoQA-SC under different bandwidth constraints. The accuracy changes of each question type relative to the upper bound are marked. In general, the accuracy of all three types of questions shows a slow decrease as the BCR decreases. One can be observed that compared to causal and temporal questions, the accuracy of descriptive questions is most affected by the reduction of BCR. This suggests that answering descriptive questions may heavily rely on the completeness of the video semantics. The experiments demonstrate that the accuracy of different types of questions are influenced differently by changes of transmission bandwidth. Overall, although we do not consider the specific designs for different question types, VideoQA-SC still achieves a good trade-off between bandwidth usage and task performance across a wide range of question types.
V Conclusion
In this paper, we have proposed an E2E SC system, VideoQA-SC, to perform VideoQA tasks in wireless networks without relying on the reconstructed source. Taking advantage of the efficient video semantic extraction and the learning-based bandwidth-adaptive DJSCC transmission, VideoQA-SC is able to fully leverage video semantic information to improve system performance for VideoQA tasks with high bandwidth-efficiency and noise-robustness. VideoQA-SC is trained in an E2E manner with the goal of maximize the VideoQA performance under bandwidth constraints. Experiments show that the proposed VideoQA-SC outperforms SSCC-based traditional systems and advanced DJSCC-based SC systems under a wide range of channel conditions and bandwidth constraints. In particular, VideoQA-SC improves $5.17\%$ VideoQA accuracy while achieves nearly $99.5\%$ bandwidth savings compared with the DJSCC-based SC system over AWGN channel when SNR is 00 dB, which demonstrates the great potential of SC system design for video applications. In this paper, we do not consider the specific designs for handling different question types, the accuracy of which are influenced differently by changes of transmission bandwidth. The future work will focus on the utilization of question types in semantic extraction and coding to make better trade-off between bandwidth usage and task performance across a wide range of question types.